# Can Direct Latent Model Learning Solve Linear Quadratic Gaussian Control?

### v2

@@ Appendix C Certainty equivalent linear quadratic control
(Mania et al., 2019) has studied this problem in the infinite-horizon LTI setting; here we extend their result to the finite-horizon LTV setting.
We adopt the following compact formulation of a finite-horizon LTV system, introduced in (Zhang et al., 2021, Section 3), to reduce our setting to the infinite-horizon LTI setting; since noisy and noiseless LQR has the same associated Riccati equation and optimal controller, we consider the noiseless case here:
$A^{\ast}=\begin{bmatrix}0_{d_{x}\times d_{x}T}&0_{d_{x}\times d_{x}}\\ \textnormal{diag}(A^{\ast}_{0},\ldots,A^{\ast}_{T-1})&0_{d_{x}T\times d_{x}}\end{bmatrix},\quad B^{\ast}=\begin{bmatrix}0_{d_{x}\times d_{u}T}\\ \textnormal{diag}(B^{\ast}_{0},\ldots,B^{\ast}_{T-1})\end{bmatrix},$
$Q^{\ast}=\textnormal{diag}(Q^{\ast}_{0},\ldots,Q^{\ast}_{T}),\quad R^{\ast}=\textnormal{diag}(R^{\ast}_{0},\ldots,R^{\ast}_{T-1}),\quad K=[\textnormal{diag}(K_{0},\ldots,K_{T-1}),0_{d_{u}T\times d_{x}}].$
The control inputs using state feedback controller $(K_{t})_{t=0}^{T-1}$ can be characterized by $u=Kx$. Let $(P_{t}^{K})_{t=0}^{T}$ be the associated cumulative cost matrix starting from step $t$. Then
and $P^{K}:=\textnormal{diag}(P_{0}^{K},\ldots,P_{T}^{K})$ is the solution to
Similarly, the optimal cumulative cost matrix
$P^{\ast}:=\textnormal{diag}(P^{\ast}_{0},\ldots,P^{\ast}_{T}),$
produced by the RDE (2.4) in system $(A^{\ast},B^{\ast},R^{\ast},Q^{\ast})$ (that is, system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$), satisfies
Let $\hat{P}=\textnormal{diag}(\hat{P}_{0},\ldots,\hat{P}_{T})$ be the optimal cumulative cost matrices in system $(\hat{A},\hat{B},\hat{Q},R^{\ast})$ (that is, system $((\hat{A}_{t},\hat{B}_{t},R^{\ast}_{t})_{t=0}^{T-1},(\hat{Q}_{t})_{t=0}^{T})$) by the RDE (2.4). Define $K^{\ast}:=[\textnormal{diag}(K^{\ast}_{0},\ldots,K^{\ast}_{T}),0_{d_{u}T\times d_{x}}]$, where $(K^{\ast}_{t})_{t=0}^{T-1}$ is the optimal controller in system $(A^{\ast},B^{\ast},Q^{\ast},R^{\ast})$, and define $\hat{K}$ similarly for system $(\hat{A},\hat{B},\hat{Q},R^{\ast})$. By definition, $(A^{\ast},B^{\ast},Q^{\ast})$ is stabilizable and observable in the sense of LTI systems. Therefore, by (Mania et al., 2019, Propositions 1 and 2), there exists dimension-free constant $\epsilon_{0}>0$ with $\epsilon_{0}^{-1}$ depending polynomially on system parameters such that as long as $\epsilon\leq\epsilon_{0}$,
$\|\hat{P}-P^{\ast}\|_{2}=\mathcal{O}(\epsilon),\quad\|\hat{K}-K^{\ast}\|_{2}=\mathcal{O}(\epsilon),$
and that $\hat{K}$ stabilizes system $(A^{\ast},B^{\ast})$. By (Fazel et al., 2018, Lemma 12),
$J((\hat{K}_{t})_{t=0}^{T-1})-J((K^{\ast})_{t=0}^{T-1})=\;$ $\sum\nolimits_{t=0}^{T}\mathrm{tr}(\Sigma_{t}(\hat{K}_{t}-K^{\ast}_{t})^{\top}(R^{\ast}_{t}+(B^{\ast}_{t})^{\top}P^{\ast}_{t+1}B^{\ast}_{t})(\hat{K}_{t}-K^{\ast}_{t}))$
=\;= $\mathrm{tr}(\Sigma(\hat{K}-K^{\ast})^{\top}(R^{\ast}+(B^{\ast})^{\top}P^{\ast}B^{\ast})(\hat{K}-K^{\ast})),$
where $\Sigma=\textnormal{diag}(\Sigma_{0},\ldots,\Sigma_{T})$ and $\Sigma_{t}$ is $\mathbb{E}[x_{t}x_{t}^{\top}]$ in system $(A^{\ast},B^{\ast})$ under state feedback controller $\hat{K}$. As a result,
$J((\hat{K}_{t})_{t=0}^{T-1})-J((K^{\ast})_{t=0}^{T-1})\leq\;$ $\|\Sigma\|_{2}\|R^{\ast}+(B^{\ast})^{\top}P^{\ast}B^{\ast}\|_{2}\|\hat{K}-K^{\ast}\|_{F}^{2}.$
we conclude that
$J((\hat{K}_{t})_{t=0}^{T-1})-J((K^{\ast})_{t=0}^{T-1})=\mathcal{O}((d_{x}\wedge d_{u})T\epsilon^{2}).$

### v3

@@ Acknowledgment
Since $x=[I_{d},I_{d}][x^{\ast};\delta]$, we have
$\mathbb{E}[xx^{\top}]=\;$ $[I_{d},I_{d}]\begin{bmatrix}\mathbb{E}[x^{\ast}(x^{\ast})^{\top}]&\mathbb{E}[x^{\ast}\delta^{\top}]\\ \mathbb{E}[\delta(x^{\ast})^{\top}]&\mathbb{E}[\delta\delta^{\top}]\end{bmatrix}\begin{bmatrix}I_{d}\\ I_{d}\end{bmatrix}$=\; $\mathbb{E}[x^{\ast}(x^{\ast})^{\top}]+\mathbb{E}[\delta\delta^{\top}]+\mathbb{E}[x^{\ast}\delta^{\top}]+\mathbb{E}[\delta(x^{\ast})^{\top}].$
Hence, by Lemma 13,
$\|\mathbb{E}[xx^{\top}]-\mathbb{E}[x^{\ast}(x^{\ast})^{\top}]\|_{2}\leq\|\mathbb{E}[\delta\delta^{\top}]\|_{2}+2\|\mathbb{E}[x^{\ast}(x^{\ast})^{\top}]\|^{1/2}_{2}\|\mathbb{E}[\delta\delta^{\top}]\|^{1/2}_{2},$
which completes the proof. ∎
Lemma 15.
Let $(x_{i})_{i=1}^{n}$ be $d$-dimensional random vectors defined on the same probability space. Then,
$\Big{\|}\mathbb{E}\Big{[}\Big{(}\sum\nolimits_{i=1}^{n}x_{i}\Big{)}\Big{(}\sum\nolimits_{i=1}^{n}x_{i}\Big{)}^{\top}\Big{]}\Big{\|}_{2}^{1/2}\leq\sum\nolimits_{i=1}^{n}\|\mathbb{E}[x_{i}x_{i}^{\top}]\|_{2}^{1/2}.$
Proof.
By the triangle inequality,
$\Big{\|}\mathbb{E}\Big{[}\Big{(}\sum\nolimits_{i=1}^{n}x_{i}\Big{)}\Big{(}\sum\nolimits_{i=1}^{n}x_{i}\Big{)}^{\top}\Big{]}\Big{\|}_{2}=\Big{\|}\mathbb{E}\Big{[}\sum\nolimits_{1\leq i,j\leq n}x_{i}x_{j}^{\top}\Big{]}\Big{\|}_{2}\leq\sum\nolimits_{1\leq i,j\leq n}\|\mathbb{E}[x_{i}x_{j}^{\top}]\|_{2}.$
By Lemma 13, for any $1\leq i,j\leq n$,
$\|\mathbb{E}[x_{i}x_{j}^{\top}]\|_{2}\leq\|\mathbb{E}[x_{i}x_{i}^{\top}]\|_{2}^{1/2}\cdot\|\mathbb{E}[x_{j}x_{j}^{\top}]\|_{2}^{1/2}.$
Hence,
$\Big{\|}\mathbb{E}\Big{[}\Big{(}\sum\nolimits_{i=1}^{n}x_{i}\Big{)}\Big{(}\sum\nolimits_{i=1}^{n}x_{i}\Big{)}^{\top}\Big{]}\Big{\|}_{2}\leq\;$ $\sum\nolimits_{1\leq i,j\leq n}\|\mathbb{E}[x_{i}x_{i}^{\top}]\|_{2}^{1/2}\cdot\|\mathbb{E}[x_{j}x_{j}^{\top}]\|_{2}^{1/2}$
$=\;$ $\Big{(}\sum\nolimits_{i=1}^{n}\|\mathbb{E}[x_{i}x_{i}^{\top}]\|_{2}^{1/2}\Big{)}^{2},$
which completes the proof. ∎